OpenStudy (loser66):

Please, help int (cl A) = int A? I would like to know is there any way to figure out the answer before seeking a counterexample.

OpenStudy (loser66):

@wio

OpenStudy (loser66):

For example: if the problem is : \( cl A\cap A = A\) the answer is yes, since cl A = A \(\cup \) bd A, hence \(A \subset \) cl A, which gives us the expression.

OpenStudy (anonymous):

I don't know what \(cl\) is.

OpenStudy (zzr0ck3r):

closer of A

OpenStudy (loser66):

to the problem above, I have to use try and error to find out the answer is NO. And we know that sometimes some special cases are true and we are lucky enough to get those true but actually, it is wrong.

OpenStudy (loser66):

closure of A

OpenStudy (loser66):

closure of A = A \(\cup\) boundary of A

OpenStudy (zzr0ck3r):

Its seems true to me. I cant think f any C.E.'s

OpenStudy (anonymous):

I don't remember what the closure of a set is.

OpenStudy (anonymous):

Oh.

OpenStudy (loser66):

@zzr0ck3r it is false

OpenStudy (zzr0ck3r):

Ive seen crazier things:)

OpenStudy (anonymous):

I think you should just consider two cases: 1) A is closed. 2) A is not closed, and the closure had additional elements.

OpenStudy (loser66):

If A ={[-1,1]\{0} in R} then bd A = {-1,0,1} and cl A = A or {-1,0,1} when int A = (-1,1) without 0

OpenStudy (zzr0ck3r):

AH so its true if A is open?

OpenStudy (loser66):

if it is true, it must be true for all.

OpenStudy (anonymous):

Okay, how about this...

OpenStudy (anonymous):

\[ int(A) = A-bd(A) \]And so \[ int(cl(A)) = cl(A) - bd(cl(A)) \]

OpenStudy (anonymous):

\[ [A\cup bd(A) ] \setminus bd(cl(A)) \]

OpenStudy (loser66):

then?

OpenStudy (anonymous):

\[ bd(cl(A)) = bd(A\cup bd(A)) \]

OpenStudy (anonymous):

Replies are out of order?!

OpenStudy (anonymous):

Is there any properties for the border of the union of two sets?

OpenStudy (loser66):

nope

OpenStudy (anonymous):

Really?

OpenStudy (anonymous):

Well, then I'm kinda stuck.

OpenStudy (loser66):

oh, is it valid? \(int \color{red}{(cl A)}= int \color{red}{A}\) which gives us cl A = A and cl A = A iff A is closed, so that the expression is wrong.

OpenStudy (anonymous):

That would only be valid if the inverse operation of \(int\) is a function.

OpenStudy (loser66):

I "cut" some arguments, so that it is not neatly at all. But the idea is that, suppose the expression is true, then lead to contradiction

OpenStudy (anonymous):

If the inverse operation of \(int\) results in two outputs, then that won't work.

OpenStudy (loser66):

hey, above, you use exactly that method, I mean you just replace A by cl A

OpenStudy (anonymous):

No, I was just trying to manipulate the expressions.

OpenStudy (loser66):

:)

OpenStudy (anonymous):

I didn't even state whether or not they were equivalent.

OpenStudy (loser66):

16 comment box above, you said int(A)=A−bd(A) And so int(cl(A))=cl(A)−bd(cl(A)) Is it not that just replace A by cl A to go from the first line to the second line?

OpenStudy (anonymous):

I was doing replacement.

OpenStudy (anonymous):

That isn't the same as an inverse.

OpenStudy (anonymous):

Okay, so suppose \[ f(int(x)) = x \]

OpenStudy (anonymous):

Then \(f^{-1}(x) = int(x)\).

OpenStudy (anonymous):

But is \(f\) actually a function?

OpenStudy (loser66):

no, it is not. it is a set

OpenStudy (anonymous):

\[ f((0,1)) =[0,1] \]and \[ f((0,1)) = (0,1) \]So therefore it isn't a function. That means that we can't apply \(f\) to both sides.

OpenStudy (loser66):

I got you, :)

OpenStudy (anonymous):

Wait a moment... can we say anything about the interior of a union of sets?

OpenStudy (anonymous):

like \(int(A\cap B)\)?

OpenStudy (loser66):

We have some properties: Union of an arbitrary collection of open subsets of M is open Union of a finite number of closed subsets of M is closed

OpenStudy (loser66):

so, if A, B closed then (A intersection B) is closed

OpenStudy (loser66):

because A , B is finite number of set.

OpenStudy (anonymous):

The boundary is always a closed set?

OpenStudy (loser66):

not always, we have some set called "clopen", it is not open nor closed, so that the boundary is the same.

OpenStudy (anonymous):

How can it be clopen?

OpenStudy (loser66):